# The Empirical Impact of Neural Parameter Symmetries, or Lack Thereof

### v2

@@ Abstract
Many algorithms and observed phenomena in deep learning appear to be affected by parameter symmetries — transformations of neural network parameters that do not change the underlying neural network function. These include linear mode connectivity, model merging, Bayesian neural network inference, metanetworks, and several other characteristics of optimization or loss-landscapes. However, theoretical analysis of the relationship between parameter space symmetries and these phenomena is difficult. In this work, we empirically investigate the impact of neural parameter symmetries by introducing new neural network architectures that have reduced parameter space symmetries. We develop two methods, with some provable guarantees, of modifying standard neural networks to reduce parameter space symmetries. With these new methods, we conduct a comprehensive experimental study consisting of multiple tasks aimed at assessing the effect of removing parameter symmetries. Our experiments reveal several interesting observations on the empirical impact of parameter symmetries; for instance, we observe linear mode connectivity between our networks without alignment of weight spaces, and we find that our networks allow for faster and more effective Bayesian neural network training.
@@ 1 Introduction
These two approaches are inspired by previous work, which shows that both symmetries of computation graphs [38] and equivariances of nonlinearities [19] induce parameter symmetries in standard neural networks. We theoretically prove that both of our approaches remove parameter symmetries under certain conditions. Our Asymmetric networks are similar structurally to standard networks and can be trained with standard backpropagation and first-order optimization algorithms like Adam. Thus, they are a reasonable “counterfactual” system for studying neural networks without parameter symmetries.
@@ 3 Related Work
Constraints and post-processing to break parameter space symmetries. Several works develop methods for constraining or post-processing the weights of a single neural network to remove ambiguities from parameter space symmetries. This includes methods to remove scaling symmetries induced by normalization layers or positively-homogeneous nonlinearities such as $\mathrm{ReLU}$ [5, 54, 53, 35], methods to remove permutation symmetries [54, 53, 70, 35], and methods to remove sign symmetries induced by odd activation functions [70]. Unlike these previous works, we develop neural network architectures that have reduced parameter space symmetries. Our models are optimized using standard unconstrained gradient-descent based methods like Adam. Hence, our networks do not require any non-standard optimization algorithms such as manifold optimization or projected gradient descent [5, 54], nor do they require post-training-processing to remove symmetries or special care during analysis of parameters (such as geodesic interpolation in a Riemannian weight space [53]).
@@ 4 Asymmetric Networks
We develop two methods of parameterizing neural network architectures without parameter symmetries, both of which are justified by theoretical results. We first focus on the case of fully-connected MLPs with no biases, which take the form $f_{\theta}(x)=\mathbf{W}_{L}\sigma(\mathbf{W}_{L-1}\cdots\sigma(\mathbf{W}_{1}x))$ for weights $\theta=(\mathbf{W}_{L},\ldots,\mathbf{W}_{1})$ and nonlinearity $\sigma$. Then in Section 4.3, we discuss how we use these approaches to remove parameter symmetries in other architectures as well.
@@ 4.1 Computation Graph Approach ($\mathbf{W}$-Asymmetric Networks)
In practice, we generate a binary mask $M$ by randomly selecting a subset of $n_{\mathrm{fix}}$ fixed elements for each row. For the fixed entries, we sample them from a normal distribution $\mathcal{N}(0,\kappa I)$ with standard deviation $\kappa>0$ that we tune. Our asymmetric linear layer can be written as
@@ 4.2.1 FiGLU: the Fixed Gated Linear Unit Nonlinearity
for a randomly sampled, untrained matrix $\mathbf{{\color[rgb]{.75,0,.25}F}}$. In this work, we sample $\mathbf{{\color[rgb]{.75,0,.25}F}}$ as an i.i.d Gaussian matrix with variance that we tune. This nonlinearity is similar to Swish / SiLU [56, 25] with an additional matrix $\mathbf{{\color[rgb]{.75,0,.25}F}}$ to mix feature dimensions (to break permutation equivariance), and it is also similar to a gated linear unit (GLU) with no trainable parameters [8]. Thus, we call our nonlinearity FiGLU: the Fixed Gated Linear Unit.
@@ 5.1 Linear Mode Connectivity without Permutation Alignment
Experimental Setup. We consider several networks and tasks: MLPs on MNIST, ResNets [22] on CIFAR-10, and Graph Neural Networks [72] on ogbn-arXiv [26]. For each architecture and task, we compute the midpoint test loss barrier: $L(\frac{1}{2}\theta_{1}+\frac{1}{2}\theta_{2})-\frac{1}{2}(L(\theta_{1})+L(\theta_{2}))$. This measures how much worse the interpolated network with parameters $\frac{1}{2}\theta_{1}+\frac{1}{2}\theta_{2}$ is than the original networks with parameters $\theta_{1}$ and $\theta_{2}$. We measure this barrier for standard networks, pairs of networks aligned with Git-Rebasin [1], and networks with our two approaches ($\sigma$-Asym and $\mathbf{W}$-Asym) applied.
@@ 5.2 Bayesian Neural Networks
Table 2: Bayesian neural network results. All results (except for last column) are after 50 epochs of training. $\mathbf{W}$-Asymmetric networks tend to improve over their standard counterparts, especially early in training. 16-layer MLPs fail to train, but 16-layer $\mathbf{W}$-Asymmetric MLPs successfully train. Standard or Asymmetric networks better than their counterpart by a standard deviation are bolded.
@@ 5.3 Metanetworks
Experimental setup. We experiment with metanetworks on the task of predicting the CIFAR-10 test accuracy of an input image classifier, which many metanetworks have been tested on [64, 11, 80, 38]. We use metanetworks based on simple MLPs, 1D-CNN metanetworks [11], and metanetworks that are exactly invariant to permutation parameter symmetries: DeepSets [76] and StatNN [64]. We train two separate datasets of 10,000 image classifiers: one dataset of small ResNet models, and one dataset of $\mathbf{W}$-Asymmetric ResNet models. More information on the data, metanetworks, and training details are in Appendix E.3.
@@ 6 Discussion
While many properties of Asymmetric networks are in line with our hypotheses and intuition about the impact of removing parameter symmetries, there are many unexpected effects and unanswered questions that are promising to further investigate. For instance, we did not extensively explore Asymmetric networks in the context of model interpretability, generalization measures in weight spaces, or optimization improvements, all of which are known to be influenced to some extent by parameter symmetries. Further studying the properties in Section 5.5, the dependence of behavior on the choices of Asymmetry-inducing hyperparameters, and other design choices in making networks asymmetric could also bring more insights into parameter space symmetries. We believe that future theoretical and empirical study of Asymmetric networks could garner many insights into the role of parameter symmetries in deep learning.
@@ Appendix D Broader Impacts
Appendix E Experimental Details
E.1 Linear Mode Connectivity Experimental Details
E.1.1 Image Classifier Interpolation
@@
E.1.2 Graph Neural Network Interpolation
@@
E.2 Bayesian Neural Network Experimental Details
@@
E.3 Metanetwork Experimental Details
E.3.1 Dataset Details
@@ E.3.2 Metanetwork Details
E.4 Monotonic Linear Interpolation Experimental Details
@@
E.5 Miscellaneous Experimental Details
@@
Appendix F arXiv Changelog

### v3

@@ Abstract
Many algorithms and observed phenomena in deep learning appear to be affected by parameter symmetries — transformations of neural network parameters that do not change the underlying neural network function. These include linear mode connectivity, model merging, Bayesian neural network inference, metanetworks, and several other characteristics of optimization or loss-landscapes. However, theoretical analysis of the relationship between parameter space symmetries and these phenomena is difficult. In this work, we empirically investigate the impact of neural parameter symmetries by introducing new neural network architectures that have reduced parameter space symmetries. We develop two methods, with some provable guarantees, of modifying standard neural networks to reduce parameter space symmetries. With these new methods, we conduct a comprehensive experimental study consisting of multiple tasks aimed at assessing the effect of removing parameter symmetries. Our experiments reveal several interesting observations on the empirical impact of parameter symmetries; for instance, we observe linear mode connectivity between our networks without alignment of weight spaces, and we find that our networks allow for faster and more effective Bayesian neural network training. Our code is available at https://github.com/cptq/asymmetric-networks.
@@ 1 Introduction
These two approaches are inspired by previous work, which shows that both symmetries of computation graphs [39] and equivariances of nonlinearities [20] induce parameter symmetries in standard neural networks. We theoretically prove that both of our approaches remove parameter symmetries under certain conditions. Our Asymmetric networks are similar structurally to standard networks and can be trained with standard backpropagation and first-order optimization algorithms like Adam. Thus, they are a reasonable “counterfactual” system for studying neural networks that are similar to standard neural networks, but that do not have as many parameter symmetries.
@@ 3 Related Work
Constraints and post-processing to break parameter space symmetries. Several works develop methods for constraining or post-processing the weights of a single neural network to remove ambiguities from parameter space symmetries. This includes methods to remove scaling symmetries induced by normalization layers or positively-homogeneous nonlinearities such as $\mathrm{ReLU}$ [6, 55, 54, 36], methods to remove permutation symmetries [55, 54, 71, 36], and methods to remove sign symmetries induced by odd activation functions [71].
Unlike these previous works, we develop neural network architectures that have reduced parameter space symmetries. Our models are optimized using standard unconstrained gradient-descent based methods like Adam. Hence, our networks do not require any non-standard optimization algorithms such as manifold optimization or projected gradient descent [6, 55], nor do they require post-training-processing to remove symmetries or special care during analysis of parameters (such as geodesic interpolation in a Riemannian weight space [54]). These methods based on constraining weights or post-processing have significantly different optimization and loss landscape properties (for instance, linear interpolation is not even well-defined on general nonlinear parameter manifolds), so they are less suitable than our Asymmetric networks for studying phenomena that may generalize to standard neural networks.
@@ 4 Asymmetric Networks
We develop two methods of parameterizing neural network architectures without parameter symmetries, both of which are justified by theoretical results. We can prove that our $\sigma$-Asym networks have permutation and scale symmetries removed, and that our $\mathbf{W}$-Asym networks have permutation symmetries removed. Although we have not formally proven that $\mathbf{W}$-Asym networks have scale symmetries removed, we believe that they do (intuitively, the fixed weights fix a scale).
We first focus on the case of fully-connected MLPs with no biases, which take the form $f_{\theta}(x)=\mathbf{W}_{L}\sigma(\mathbf{W}_{L-1}\cdots\sigma(\mathbf{W}_{1}x))$ for weights $\theta=(\mathbf{W}_{L},\ldots,\mathbf{W}_{1})$ and nonlinearity $\sigma$. Then in Section 4.3, we discuss how we use these approaches to remove parameter symmetries in other architectures (e.g. CNNs, GNNs) as well.
@@ 4.1 Computation Graph Approach ($\mathbf{W}$-Asymmetric Networks)
In practice, we generate a binary mask $M$ by randomly selecting a subset of $n_{\mathrm{fix}}$ fixed elements for each row. For the fixed entries, we sample them from a normal distribution $\mathcal{N}(0,\kappa I)$ with standard deviation $\kappa>0$ that is a hyperparameter that we tune. Our asymmetric linear layer can be written as
@@ 4.2.1 FiGLU: the Fixed Gated Linear Unit Nonlinearity
for a randomly sampled, untrained matrix $\mathbf{{\color[rgb]{.75,0,.25}F}}$. Similarly to $\mathbf{W}$-Asym nets, we sample $\mathbf{{\color[rgb]{.75,0,.25}F}}$ as an i.i.d Gaussian matrix with variance that we tune. This nonlinearity is similar to Swish / SiLU [57, 26] with an additional matrix $\mathbf{{\color[rgb]{.75,0,.25}F}}$ to mix feature dimensions (to break permutation equivariance), and it is also similar to a gated linear unit (GLU) with no trainable parameters [9]. Thus, we call our nonlinearity FiGLU: the Fixed Gated Linear Unit.
@@ 4.4 Universal Approximation
We have not been able to prove a similar result for $\sigma$-Asymmetric networks. Classical universal approximation results for standard neural networks do not apply to $\sigma$-Asym nets, as they tend to assume elementwise nonlinearities.
@@ 5.1 Linear Mode Connectivity without Permutation Alignment
Experimental Setup. We consider several networks and tasks: MLPs on MNIST, ResNets [23] on CIFAR-10, and Graph Neural Networks [73] on ogbn-arXiv [27]. For each architecture and task, we compute the midpoint test loss barrier: $L(\frac{1}{2}\theta_{1}+\frac{1}{2}\theta_{2})-\frac{1}{2}(L(\theta_{1})+L(\theta_{2}))$. This measures how much worse the interpolated network with parameters $\frac{1}{2}\theta_{1}+\frac{1}{2}\theta_{2}$ is than the original networks with parameters $\theta_{1}$ and $\theta_{2}$. We measure this barrier for standard networks, pairs of networks aligned with Git-Rebasin [1], and networks with our two approaches ($\sigma$-Asym and $\mathbf{W}$-Asym) applied. To be clear, whenever we interpolate between the weights of two Asymmetric networks, they have the same exact fixed weights $\mathbf{{\color[rgb]{.75,0,.25}F}}$ (for both $\mathbf{W}$-Asym and $\sigma$-Asym) and the same exact mask $M$ (for $\mathbf{W}$-Asym).
@@ 5.2 Bayesian Neural Networks
Table 2: Bayesian neural network results. Reported loss is the negative log likelihood loss. All results (except for last column) are after 50 epochs of training. $\mathbf{W}$-Asymmetric networks tend to improve over their standard counterparts, especially early in training. 16-layer MLPs fail to train, but 16-layer $\mathbf{W}$-Asymmetric MLPs successfully train. Standard or Asymmetric networks better than their counterpart by a standard deviation are bolded.
@@ 5.3 Metanetworks
Experimental setup. We experiment with metanetworks on the task of predicting the CIFAR-10 test accuracy of an input image classifier, which many metanetworks have been tested on [65, 12, 81, 39]. We use metanetworks based on simple MLPs, 1D-CNN metanetworks [12], and metanetworks that are exactly invariant to permutation parameter symmetries: DeepSets [77] and StatNN [65]. We train two separate datasets of 10,000 image classifiers: one dataset of small ResNet models, and one dataset of $\mathbf{W}$-Asymmetric ResNet models. More information on the data, metanetworks, and training details are in Appendix F.3.
@@ 6 Discussion
While many properties of Asymmetric networks are in line with our hypotheses and intuition about the impact of removing parameter symmetries, there are many unexpected effects and unanswered questions that are promising to further investigate. For instance, we did not extensively explore Asymmetric networks in the context of model interpretability, generalization measures in weight spaces, or optimization improvements, all of which are known to be influenced to some extent by parameter symmetries. Further studying the properties in Section 5.5, the dependence of behavior on the choices of Asymmetry-inducing hyperparameters, and other design choices in making networks asymmetric could also bring more insights into parameter space symmetries.
Also, it is interesting that our $\sigma$-Asymmetric networks do not appear to break parameter symmetries as well as our $\mathbf{W}$-Asymmetric networks. We have run preliminary empirical tests on several variants of $\sigma$-Asym networks, such as: $\sigma\circ\mathbf{{\color[rgb]{.75,0,.25}F}}\circ\sigma$ as the nonlinearity, sparsifying $\mathbf{{\color[rgb]{.75,0,.25}F}}$, adding instead of multiplying the gate, using cosine instead of sigmoid, squaring instead of using sigmoid, and putting a LayerNorm in the nonlinearity. However, none of these approaches worked well. We believe that such failures may be because $\sigma$-Asymmetry breaks symmetries at the activation / neuron level, whereas $\mathbf{W}$-Asymmetry breaks symmetries in the larger space of weights (for more evidence, see Appendix E, where we show that fixing biases at neurons also fails to effectively remove symmetries). These curiosities provide interesting directions for future work.
All in all, we believe that future theoretical and empirical study of Asymmetric networks could garner many insights into the role of parameter symmetries in deep learning.
@@ Appendix D Broader Impacts
Appendix E Experimental Ablations
Here, we present experiments with various ablations. These were in-part suggested by anonymous reviewers for the NeurIPS 2024 conference (thanks!).
E.1 Matching Learnable Parameters
Table 5: Number of learnable parameters of Standard / $\sigma$-Asym and $\mathbf{W}$-Asym nets for our experiments.
Experiment / Architecture Standard / $\sigma$-Asym W-Asym
5.1 MLP 935,434 834,570
5.1 ResNet 1x 272,474 230,024
5.1 ResNet 8x 17,289,866 16,273,946
5.1 GNN 176,424 171,576
5.2 MLP-8 3,242,146 3,324,466
5.2 MLP-16 5,960,242 5,796,002
5.2 ResNet-20 1x 1,356,098 1,143,858
5.2 ResNet-20 2x 5,410,386 5,044,756
5.2 ResNet-110 1x 8,620,418 7,371,378
5.2 ResNet-110 2x 34,512,276 32,014,996
5.2 ResNet-20 2x 5,410,386 5,044,756
5.3 ResNet 78,042 60,634
5.4 MLI ResNet 78,042 60,634
Table 6: Metanetwork performance, including results with smaller standard networks (Smaller ResNet) at same number of parameters at $\mathbf{W}$-Asym. There is no substantial difference — $\mathbf{W}$-Asym ResNets are still substantially easier to predict performance of.
ResNet Smaller ResNet $\mathbf{W}$-Asym ResNet
$R^{2}$ $\tau$ $R^{2}$ $\tau$ $R^{2}$ $\tau$
[…]
Table 7: Bayesian NN test accuracy after 25 epochs. Decreasing standard ResNet20 parameters to match that of $\mathbf{W}$-Asym ResNet20 does not substantially change performance.
Base Network Test Accuracy
$\mathbf{W}$-Asym ResNet20 $\mathbf{49.3}\pm 0.4$
Standard ResNet20 $46.8\pm 0.9$
Smaller Standard ResNet20 $46.5\pm 1.1$
For the experiments in the main text, our $\mathbf{W}$-Asym networks often had less learnable parameters than the standard networks that they were compared against (since we take them to have the same architecture, but the $\mathbf{W}$-Asym approach fixes certain learnable parameters to constants). See Table 5 for the learnable parameter counts.
In this section, we control for this, by decreasing the width of the standard networks so that they have the same number of parameters as the $\mathbf{W}$-Asymmetric networks they are compared against. Results are shown in Tables 6 and 7. We see that there is little to no change in metanetwork performance or Bayesian neural network performance for these smaller standard networks, so our results from the main text all still hold.
E.2 Changing Number of Warmup Steps
Table 8: Test loss barrier when changing warmup steps. Results are very similar when lowering number of warmup epochs ($\mathbf{W}$-Asym interpolates significantly better than Git-ReBasin). Adam optimizer with learning rate 1e-2 (ResNet20) and 1e-3 (GNN) is used.
ResNet20 (ReBasin) $\mathbf{W}$-Asym ResNet20 GNN (ReBasin) $\mathbf{W}$-Asym GNN
[…]
Altintas et al. [3] showed that amount of learning rate warmup can affect the extent of linear mode connectivity between training runs. In Table 8, we see that varying the number of warmup epochs between 1 and 20 does not change the qualitative results much on test loss barrier for our linear mode connectivity experiments.
E.3 Failure to Break Symmetries by Fixing Biases
Table 9: Loss barrier when fixing biases to attempt to break symmetries for standard ResNet20 on CIFAR-10. Biases are between $[-k,k]$. Barriers are much larger than for $\mathbf{W}$-Asymmetric Nets (.$934\pm.72$) and $\sigma$-Asym Nets ($2.521\pm.46$).
$k$ Loss Barrier
1 $5.81\pm 3.67$
3 $3.76\pm 0.38$
9 $4.62\pm 1.03$
27 $10.6\pm 4.29$
Intuitively, one way to break permutation symmetries of an MLP is to order the hidden neurons in some way. One possible way to do this is to fix the biases (so that they are untrained). As shown in Table 9, a basic attempt at this fails, and has a much larger test barrier than our $\mathbf{W}$-Asym and $\sigma$-Asym networks.
Appendix F Experimental Details
F.1 Linear Mode Connectivity Experimental Details
F.1.1 Image Classifier Interpolation
@@
F.1.2 Graph Neural Network Interpolation
@@
F.2 Bayesian Neural Network Experimental Details
@@
F.3 Metanetwork Experimental Details
F.3.1 Dataset Details
@@ F.3.2 Metanetwork Details
F.4 Monotonic Linear Interpolation Experimental Details
@@
F.5 Miscellaneous Experimental Details
@@
Appendix G arXiv Changelog
@@
Version 3 (October 14 2024): Added link to code. Also added some additional experiments/ablations, and made various other small changes for the NeurIPS camera-ready version.